Find an equation of a polynomial function of least degree having complex zeros: -i & -2i. X-intercept: (2,0). Y-intercept: (0,-2)
Answer by stanbon(75887)

Find an equation of a polynomial function of least degree having complex zeros: -i & -2i. X-intercept: (2,0). Y-intercept: (0,-2)
-----
If I assume the coefficients of the polynomial function
are Real Numbers, then i and 2i are also zeros.
-------
f(x) = a(x-i)(x+i)(x-2i)(x+2i)(x-2)
----
To solve for "a", use the point (0,-2)
-2 = a(-i)(i)(-2i)(2i)(-2)
1 = a(1)(4)
a = 1/4
---
Equation:
f(x) = (1/4)(x^2+1)(x^2+4)(x-2)
Degree: 5
==================
Cheers,
Stan H.
=======================